RC meeting and Run Plan, Tuesday November 15th
From SBS wiki
Return to the RC Meeting Calendar
Contents
RC Meeting Information
Time:at 3:30pm today
- Location: Counting house conference room
- Join ZoomGov Meeting
- https://jlab-org.zoomgov.com/j/1619032184?pwd=aWhTVWxya1ZZQzA1K2ZPSTd2ZX...
- Meeting ID: 161 903 2184
- Passcode: 260715
- Dial by your location
- +1 669 254 5252 US (San Jose)
- +1 646 828 7666 US (New York)
- +1 551 285 1373 US
- +1 669 216 1590 US (San Jose)
- 833 568 8864 US Toll-free
- Find your local number: https://jlab-org.zoomgov.com/u/acj3pTicMx
Useful Links
Current Status
Accelerator and Hall Status
- Beam maintenance started at 6:45am
- Hall is in "Controlled Access". Attached is a list of work requests at the end of this report.
- We took production data at 45muA all of swing and owl shifts yesterday.
- Target tests were done before sweeping the hall this morning.
- As soon as we lock up the hall, target experts will perform additional tests.
- Beam is expected to be back around 8pm.
- MCC is happy to support us for mini spin dance measurement in association with the Moller measurement on Thursday.
- Won't do BCM calibration tomorrow.
Current Activity and Issues
DAQ
- DAQ gets stuck somewhat frequently without obvious warning/error message. Shift crew gets confused and it may end up taking hours to fix. Is it possible to force CODA to throw better error messages? - Alex
Beamline
- Dump right Ion chamber unresponsive. For now director of operations agreed to allow running with just one dump IC. - J. Kowal and team will work on it (& Compton IC) tomorrow
Run Plan
Current Plan
- Perform target tests
- Do beam checkout
- Continue taking production data
- 43uA of beam current | Use ps3=ps5=0, ps1=9, ps2=15, rest =-1 | One hour long runs
- Move the target to the pickup coil position and do an NMR sweep every 2/3 hours
- Return to step 1.
Requests
- Moller measurement - Thursday day shift (Mini Spin dance?)
- Optics runs with sieve in
- Calibration runs on H2
- GEM luminosity study
Subsystems Updates
Polarized Target
- Polarization: 42.4% (12:17pm)
Quick Summary on the performance of every sub-system during GEn
GRINCH
BBGEM
SBSGEM
Instructions for Shift Crew
- Keep an eye out for drifting BPM positions and TS1-BBHI rate. Post a screenshot of the relevant strip-charts and GUI twice every shift.
- If beam has been down due to accelerator issues perform a checkout of the beam profile by doing harp scans and check beam centering on target using the Carbon hole target, see Simona's procedure: procedure, If in doubt about whether you should go through the beam checkout procedure, call this number: 7578693131.
- Cross check the rate with what is expected: He3, H2
- Cross check with the expected ion chamber readings: log
- If holding field is changed for any reason, do NOT perform an NMR measurement.
Be sure to follow these procedures:
- Power down SBS, BB, and Detector HV before controlled access and allowing RadCon in the hall
- Communicate to the TO that the SBS and BB magnets will be powered down, so they can update the holding field
- See the beam recovery procedures at: HOW_TOs#Beam_Recovery_Procedure
- See the instructions to set the raster at: HOW_TOs#How_to_ask_MCC_to_set_the_desired_raster_size_of_6_mm
Miscellaneous
Experimental progress (for MCC planning meeting)
- 8.6% of accumulated charge for production data (19.9 Coulombs out of 232)
- 36% Experimental run time (51 out of 141 days)
- Long term "to do" items; not urgent, some can be done when 3He target not available
- Luminosity scans with GEMs. See procedure.
- Opportunistic BCM calibration (higher current) before pass change?
- INFN LV Scan and DAQ Stability - maybe 20-30 minutes if nothing else can run
- Work and testing on LV supplies was performed during opportunistic access. See log.
- Misc. GEM commissioning items: https://sbs.jlab.org/wiki/index.php/CommissioningPlan#BB.2FSBS_GEM_Commissioning_.28Continuation.29